# Jizhou Liu Booth School of Business

### v1

@@ 3.1 Overview of Conditional Randomization Tests
The classical Fisher Randomization Test simulates the sampling distribution of the test statistic according to the actual treatment variation in the experiment (Imbens and Rubin, 2015, Chapter 5). While this procedure is valid for testing the sharp null hypothesis under interference, it is generally not valid for non-sharp hypotheses, which do not specify the complete schedule of potential outcomes. The null hypotheses defined in the previous section fall into this category. For instance, if we observe outcome $Y_{ij}(1,0)$ for a treated buyer $i$ and control seller $j$, then we cannot impute the outcome $Y_{ij}(0,0)$ under $H_{0}^{seller}$—this null hypothesis is not sharp.
@@
Such a test is finite-sample valid for a non-sharp null hypothesis, provided that the conditioning event, $\mathcal{C}$, is constructed in a way such that the potential outcomes for all units and assignments in $\mathcal{C}$ can be imputed under the null hypothesis (Basse et al., 2019, Theorem 1). Sampling from (3) may be challenging in general. However, under certain conditions on the design, $P(\mathbf{W})$, and the conditioning mechanism, $P(\mathcal{C}|\mathbf{W})$, the distribution in (3) can be reduced to a permutation distribution that is easy to sample from. We will consider this approach in the following sections.
@@ 3.2 Testing spillover effects
In this section, we develop conditional randomization tests for the spillover hypotheses $H_{0}^{\text{buyer}}$ and $H_{0}^{\text{seller}}$ defined above. Under symmetric designs, our randomization procedure entails straightforward permutations of treatment assignments on one side (buyer or seller), conditioned on the assignment of the other side (seller or buyer, respectively). We will begin with the buyer spillover hypothesis and describe the associated randomization procedure in detail. The analysis of the seller spillover hypothesis will be brief as it is completely symmetrical to the buyer case.
For the buyer spillover hypothesis 2.1, we propose to condition on control sellers and fix those sellers to control across randomizations to ensure that the potential outcomes can be imputed. Moreover, we will condition on the marginal number of treated buyer-sellers observed in the sample. Under Assumption 2.2, the conditional randomization distribution is exchangeable, resulting in a permutation test that can be efficiently implemented. Formally, we define the following conditioning events:
@@
where =$\mathcal{M}=\{vu^{\top}:v\in\mathbb{W}^{B},u\in\mathbb{W}^{S}\text{ s.t. }\sum_{i=1}^{I}v_{i}=\sum_{i=1}^{I}w^{obs,B}_{i},\sum_{i=j}^{J}u_{j}=\sum_{i=j}^{J}w^{obs,S}_{j}\}$. Here, we use the superscript ‘obs’ to emphasize that we are referring to the actual treatments realized in the observed data. Thus, $\mathcal{U}$ denotes the buyer-seller pairs for which the seller part of the pair is assigned control. Set $\mathcal{W}$ denotes the assignments that these sellers in control, while $\mathcal{M}$ is the set of assignments that match the marginal number of treated buyers and sellers observed in the sample. We note that all these sets are random as they all depend on the observed treatment assignment vector, $\mathbf{W}^{obs}$. We are now ready to define the main randomization test for the buyer spillover hypothesis.
Procedure 3.1 (Test for Buyer Spillover Effect $H_{0}^{buyer}$).
@@
    where
    $n_{1}=\sum_{(i,j)\in\mathcal{U}}(1-w^{S}_{j})w^{B}_{i},\quad n_{0}=\sum_{(i,j)\in\mathcal{U}}(1-w^{S}_{j})(1-w^{B}_{i}).$
@@
    $\frac{1}{L+1}\left[1+\sum_{l=1}^{L}\mathbb{I}\left\{T(\mathbf{W}^{(l)}\mid\mathbf{Y}^{obs},\mathcal{C})>T(\mathbf{W}^{obs}\mid\mathbf{Y}^{obs},\mathcal{C})\right\}\right]\leq\alpha.$
@@
Remark 3.2.
@@
See Appendix A.1 for more details and the proof. That is, we repeatedly sample from the buyer-side design, conditioning on the treated fraction in the observed buyer-side assignment. However, sampling from such conditional distribution—e.g., through rejection sampling—may be computationally challenging depending on the particular experimental design.
@@
$\mathcal{U}=\{(i,j):w^{obs,B}_{j}=0\}\text{ and }\mathcal{W}=\{w^{obs,B}u^{\top}:u\in\mathbb{W}^{S}\}\cap\mathcal{M}.$
@@
The validity of our proposed randomization tests in Procedure 3.1 follows by adapting Theorem 1 of Basse et al. (2019) in the setting of two-sided experiments. We provide the statement of the validity result in the following theorem.
@@
    Under Assumption 2.2, the testing procedure $\phi(\mathbf{W}^{obs};\mathcal{C})$, as defined in Procedure 3.1, is valid. Specifically, for any significance level $\alpha\in(0,1)$, the expectation $E[\phi(\mathbf{W}^{obs};\mathcal{C})\mid\mathcal{C}]\leq\alpha$ holds under the null hypothesis $H_{0}^{buyer}$ (or $H_{0}^{total}$).
@@
    If Step (b) of Procedure 3.1 is replaced with the one in Remark 3.2, the validity of $\phi(\mathbf{W}^{obs};\mathcal{C})$ is maintained without assuming Assumption 2.2.
@@ 3.3.1 Testing Procedure with $k$-Block Conditioning Event
Towards testing the total effect, let $\mathcal{I}^{(k)}(\mathbf{W}^{obs})$ denote a partition of the buyer set $[I]$ into $k$ equal-sized non-overlapping subsets with all buyers having the same treatment status within each subset. [^2] We assume $I$ and $\sum_{i=1}^{I}w_{i}^{obs,B}$ are both divisible by $k$ for simplicity. When this is not the case, our methods remain valid, and the framework can be extended accordingly, though at the cost of more cumbersome notation. That is,
@@
Similarly, we define $\mathcal{J}^{(k)}(\mathbf{W}^{obs})$ as a partition of the seller set $[J]$ into $k$-sized subsets with sellers having the same treatment status within each subset. Next, define
$\mathcal{U}^{(k)}(\mathbf{W}^{obs})=\big{\{}\{(i,j):w_{i}^{obs,B}=w_{j}^{obs,S},i\in\mathcal{I}_{s},j\in\mathcal{J}_{s}\},1\leq s\leq\min(I/k,J/k)\big{\}},$ (6)
where $\mathcal{I}_{s},\mathcal{J}_{s^{\prime}}$ are defined above. That is, $\mathcal{U}^{(k)}(\mathbf{W}^{obs})$ is the collection of buyer-seller pairs constructed from the cross-product of $\mathcal{I}^{(k)}(\mathbf{W}^{obs})$ and $\mathcal{J}^{(k)}(\mathbf{W}^{obs})$, making sure that all buyers and sellers have the same treatment status in the group. Figure 2 below illustrates this construction. In the figure, the set $\mathcal{U}^{(k)}(\mathbf{W}^{obs})$ therefore corresponds to the diagonal blocks in the shaded area of the buyer-seller array. Note how these blocks are non-overlapping, they partition the sets $[I]$ and $[J]$, and, within each block, the individual treatments of every buyer and seller are identical.
@@
Procedure 3.2 (Test Total Effect $H_{0}^{total}$).
@@
    where
    $n_{1,k}=\sum_{(i,j)\in\mathcal{U}^{(k)}(\mathbf{W}^{obs})}w^{S}_{j}w^{B}_{i},\quad n_{0,k}=\sum_{(i,j)\in\mathcal{U}^{(k)}(\mathbf{W}^{obs})}(1-w^{S}_{j})(1-w^{B}_{i}).$
@@
    $w^{B,(l)}$ =$=\left\{w^{B}\in\mathbb{W}^{B}\mid\begin{cases}(w^{B}_{i})_{i\in\mathcal{I}_{s}}=(w_{i}^{B,obs})_{i\in\mathcal{I}_{\pi(s)}},&\text{if }s\leq\min(I/k,J/k),\\ (w^{B}_{i})_{i\in\mathcal{I}_{s}}=(w_{i}^{B,obs})_{i\in\mathcal{I}_{s}},&\text{otherwise}.\end{cases}\right\}$
    $w^{S,(l)}$ =$=\left\{w^{S}\in\mathbb{W}^{S}\mid\begin{cases}(w^{S}_{j})_{j\in\mathcal{J}_{s}}=(w_{j}^{B,obs})_{j\in\mathcal{J}_{\pi(s)}},&\text{if }s\leq\min(I/k,J/k),\\ (w^{B}_{j})_{i\in\mathcal{J}_{s}}=(w_{j}^{B,obs})_{j\in\mathcal{J}_{s}},&\text{otherwise}.\end{cases}\right\}$
@@
    Reject $H_{0}^{total}$, denoted by $\phi(\mathbf{W}^{obs};\mathcal{C}^{(k)})=1$, if
@@
Intuitively, to test $H_{0}^{total}$, Procedure 3.2 conditions on a $k$-block conditioning event —e.g., the $k$ diagonal blocks shown in Figure 2— and permutes the block treatments $(0,0)$, $(1,1)$ at the block level. We note that the diagonal structure in the blocks is not essential and alternative constructions can be valid as long as the blocks are non-overlapping in neither the buyer nor the seller dimension.
@@
Consider an independent two-side ranomdized design (Definition 2.1) where Assumption 2.1 holds. Let $\mathcal{C}^{(k)}$ be a $k$-block conditioning event as in Definition 3.1. Suppose also that Assumption 2.2 holds. Then, $\phi(\mathbf{W}^{obs};\mathcal{C}^{(k)})$ defined in Procedure 3.2 is finite-sample valid, such that $\mathbb{E}[\phi(\mathbf{W}^{obs};\mathcal{C}^{(k)})]\leq\alpha$ for any finite $I,J$.
@@ 3.3.2 Selecting $k$ Based on Power Considerations
$\phi_{|\mathcal{U}|,|\mathcal{W}|}\geq\frac{1}{1+Ae^{-a\tau\sqrt{C(2n/k,n/k)}}}-O\left((2nk)^{-0.5+\delta}\right)-\epsilon,$
@@
In Figure 3, we illustrate the lower bound on power as a function of the block size $k$, highlighting the trade-off between sample size and the number of unique treatment assignments as $k$ increases. While the plot identifies the optimal $k$ that maximizes power, in practice, the parameters of the lower bound function are unknown. Consequently, the optimal $k$ cannot be directly determined using Proposition 3.1. However, a heuristic approach can be used to select the largest feasible $k$ based on a predetermined maximum power threshold. Based on Proposition 3.1, $O\left((2nk)^{-0.5+\delta}\right)$ controls the maximum power of the test, while $C(2n/k,n/k)$ determines how rapidly the power function reaches its maximum (sensitivity). Therefore, to optimize the detection of treatment effects, we recommend selecting the largest feasible block size $k$ based on a predetermined maximum power. For example, setting this power to 0.95 ensures the maximal utilization of observations to detect the minimal size of the treatment effect. In practice, we can set $k\approx\frac{1}{2n(1-\beta)^{2}}$, where $\beta$ is the predetermined power level. For example, consider the completely randomized two-sided design discussed before, where choosing a block size $k=n/6$ results in $|\mathcal{W}|=C(12,6)=924$, achieving a maximum power approximately equal to 0.967.
@@ 4 Randomization Tests for Weak Null Hypotheses
In this section, we focus on testing weak null hypotheses on the treatment effects considered so far. We specifically focus on a weak null for the buyer spillover hypothesis, as the other two hypotheses of interest follow similar principles. For testing total effects, refer to the discussion in Appendix B.2.3.
Hypothesis 4.1 (Weak buyer spillover).
The weak null hypothesis of no buyer spillover effects is:
$H_{0}^{wb}:\frac{1}{IJ}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{i,j}(0,0)=\frac{1}{IJ}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{i,j}(1,0)~{}.$ (8)
There has been a growing number of works on the study of randomization tests under weak null hypotheses similar to the definition in Equation (8); see, for example, (Chung and Romano, 2013; Li and Ding, 2017; Canay et al., 2017; DiCiccio and Romano, 2017a, b; Zhao and Ding, 2021; Wu and Ding, 2021). This literature has pointed out that randomization tests based on sharp nulls are not necessarily valid under the weak null. However, properly studentized test statistics can restore the validity of sharp randomization tests even under the weak null. Following this approach, we propose a permutation test employing a studentized test statistic for testing $H_{0}^{wb}$ assuming a completely randomized design.
Procedure 4.1 (Test Weak Null $H_{0}^{wb}$).

### v2

@@ 3.1 Overview of Conditional Randomization Tests
The classical Fisher Randomization Test simulates the sampling distribution of the test statistic according to the actual treatment variation in the experiment (Imbens and Rubin, 2015, Chapter 5). While this procedure is valid for testing the sharp null hypothesis under interference, it is generally not valid for non-sharp hypotheses, which do not specify the complete schedule of potential outcomes. The null hypotheses defined in the previous section fall into this category. For instance, if we observe outcome $Y_{ij}(1,0)$ for a treated buyer $i$ and control seller $j$, then we cannot impute the outcome $Y_{ij}(0,0)$ under $H_{0}^{\text{seller}}$—this null hypothesis is not sharp in the strict sense.
@@
Such a test is finite-sample valid even for a non-sharp null hypothesis, provided that the conditioning event, $\mathcal{C}$, is constructed in a way such that the potential outcomes for all units and assignments in $\mathcal{C}$ can be imputed under the null hypothesis (Basse et al., 2019, Theorem 1). Sampling from the conditional distribution in (3) may be challenging in general. However, under certain conditions on the design, $p(\mathbf{W})$, and the conditioning mechanism, $p(\mathcal{C}|\mathbf{W})$, the distribution in (3) can be reduced to a permutation distribution that is easy to sample from. We will consider this approach in the following sections.
@@ 3.2 Testing spillover effects
In this section, we develop conditional randomization tests for the spillover hypotheses $H_{0}^{\text{buyer}}$ and $H_{0}^{\text{seller}}$. Under symmetric designs, our randomization procedure entails straightforward permutations of treatment assignments on one side (buyer or seller), conditioned on the assignment of the other side (seller or buyer, respectively). We begin with the buyer spillover hypothesis and describe the associated randomization procedure in detail. The analysis of the seller spillover hypothesis will be brief as it is completely symmetrical to the buyer case.
For the buyer spillover hypothesis 2.1, we propose to condition on control sellers and fix those sellers to control across randomizations to ensure that the potential outcomes can be imputed. Moreover, we will condition on the marginal number of treated buyer-sellers observed in the sample. Under Assumption 2.2, the conditional randomization distribution is exchangeable, resulting in a permutation test that can be efficiently implemented.
Formally, we define the following conditioning events:
@@
where $\mathcal{M}=\{vu^{\top}:v\in\mathbb{W}^{B},u\in\mathbb{W}^{S}\text{ s.t. }\sum_{i=1}^{I}v_{i}=\sum_{i=1}^{I}w^{obs,B}_{i},\sum_{j=1}^{J}u_{j}=\sum_{j=1}^{J}w^{obs,S}_{j}\}$. Here, we use the superscript ‘obs’ to emphasize that we are referring to the actual treatments realized in the observed data. Thus, $\mathcal{U}$ denotes the buyer-seller pairs for which the seller part of the pair is assigned control. Set $\mathcal{W}$ denotes the assignments that these sellers in control, while $\mathcal{M}$ is the set of assignments that match the marginal number of treated buyers and sellers observed in the sample. We note that all these sets are random as they all depend on the observed treatment assignment vector, $\mathbf{W}^{obs}$. We are now ready to define the main randomization test for the buyer spillover hypothesis.
Procedure 3.1 (Test for Buyer Spillover Effect $H_{0}^{\text{buyer}}$).
@@
    where $n_{1}=\sum_{(i,j)\in\mathcal{U}}(1-w^{S}_{j})w^{B}_{i},\quad n_{0}=\sum_{(i,j)\in\mathcal{U}}(1-w^{S}_{j})(1-w^{B}_{i}).$
@@
    $\frac{1}{L+1}\left[1+\sum_{l=1}^{L}\mathbb{I}\left\{T(\mathbf{W}^{(l)}\mid\mathbf{Y}^{obs},\mathcal{C})\geq T(\mathbf{W}^{obs}\mid\mathbf{Y}^{obs},\mathcal{C})\right\}\right]\leq\alpha.$
@@
Remark 3.2 (Randomization tests when design symmetry fails).
@@
That is, we repeatedly sample from the buyer-side design, conditioning on the treated fraction in the observed buyer-side assignment. However, sampling from such conditional distribution—e.g., through rejection sampling—may be computationally challenging depending on the particular experimental design. See Appendix A.1 for more details and the proof.
@@
$\mathcal{U}=\{(i,j):w^{obs,B}_{i}=0\}\text{ and }\mathcal{W}=\{w^{obs,B}u^{\top}:u\in\mathbb{W}^{S}\}\cap\mathcal{M}.$
@@
The validity of our proposed randomization tests in Procedure 3.1 follows by adapting Theorem 2 of Puelz et al. (2021) in the setting of two-sided experiments. We provide the statement of the validity result in the following theorem.
@@
    Under Assumption 2.2, the testing procedure $\phi(\mathbf{W}^{obs};\mathcal{C})$, as defined in Procedure 3.1, is valid. Specifically, for any significance level $\alpha\in(0,1)$, the expectation $E[\phi(\mathbf{W}^{obs};\mathcal{C})\mid\mathcal{C}]\leq\alpha$ holds under the null hypothesis $H_{0}^{\text{buyer}}$ (or $H_{0}^{\text{seller}}$).
@@
    If Step (b) of Procedure 3.1 is replaced with the one in Remark 3.2, the validity of $\phi(\mathbf{W}^{obs};\mathcal{C})$ is maintained without assuming Assumption $2.$
@@ 3.3.1 Testing Procedure with $k$-Block Conditioning Event
Towards testing the total effect, let $\mathcal{I}^{(k)}(\mathbf{W}^{obs})$ denote a random partition of the buyer set $[I]$ into $k$ equal-sized non-overlapping subsets with all buyers having the same treatment status within each subset. [^2] We assume $I$ and $\sum_{i=1}^{I}w_{i}^{obs,B}$ are both divisible by $k$ for simplicity. When this is not the case, our methods remain valid, and the framework can be extended accordingly, though at the cost of more cumbersome notation. That is,
@@
Similarly, we define $\mathcal{J}^{(k)}(\mathbf{W}^{obs})$ as a random partition of the seller set $[J]$ into $k$-sized subsets with sellers having the same treatment status within each subset. Next, define
$\mathcal{U}^{(k)}(\mathbf{W}^{obs})=\bigcup_{1\leq s\leq\min(I/k,J/k)}\{(i,j):w_{i}^{obs,B}=w_{j}^{obs,S},i\in\mathcal{I}_{s},j\in\mathcal{J}_{s}\},$ (6)
where $\mathcal{I}_{s},\mathcal{J}_{s}$ are defined above. That is, $\mathcal{U}^{(k)}(\mathbf{W}^{obs})$ is the collection of buyer-seller pairs constructed from the cross-product of $\mathcal{I}^{(k)}(\mathbf{W}^{obs})$ and $\mathcal{J}^{(k)}(\mathbf{W}^{obs})$, making sure that all buyers and sellers have the same treatment status in the group. Figure 2 below illustrates this construction. In the figure, the set $\mathcal{U}^{(k)}(\mathbf{W}^{obs})$ therefore corresponds to the diagonal blocks in the shaded area of the buyer-seller array. Note how these blocks are non-overlapping, they partition the sets $[I]$ and $[J]$, and, within each block, the individual treatments of every buyer and seller are identical.
@@
Procedure 3.2 (Test Total Effect $H_{0}^{\text{total}}$).
@@
    where $n_{1,k}=\sum_{(i,j)\in\mathcal{U}^{(k)}(\mathbf{W}^{obs})}w^{S}_{j}w^{B}_{i},\quad n_{0,k}=\sum_{(i,j)\in\mathcal{U}^{(k)}(\mathbf{W}^{obs})}(1-w^{S}_{j})(1-w^{B}_{i}).$
@@
    $w^{B,(l)}$ $:=\Bigl\{w^{B}\in\mathbb{W}^{B}\ \Big|\ (w^{B}_{i})_{i\in\mathcal{I}_{s}}=(w_{i}^{B,\mathrm{obs}})_{i\in\mathcal{I}_{\pi(s)}}\ \ \forall s\in\{1,\dots,\min(I/k,J/k)\}\Bigr\},$
    $w^{S,(l)}$ $:=\Bigl\{w^{S}\in\mathbb{W}^{S}\ \Big|\ (w^{S}_{j})_{j\in\mathcal{J}_{s}}=(w_{j}^{S,\mathrm{obs}})_{j\in\mathcal{J}_{\pi(s)}}\ \ \forall s\in\{1,\dots,\min(I/k,J/k)\}\Bigr\}.$
@@
    Reject $H_{0}^{\text{total}}$, denoted by $\phi(\mathbf{W}^{obs};\mathcal{C}^{(k)})=1$, if
@@
Intuitively, to test $H_{0}^{\text{total}}$, Procedure 3.2 conditions on a $k$-block conditioning event —e.g., the $k$ diagonal blocks shown in Figure 2— and permutes the block treatments $(0,0)$, $(1,1)$ at the block level. We note that the diagonal structure in the blocks is not essential and alternative constructions can be valid as long as the blocks are non-overlapping in neither the buyer nor the seller dimension.
@@
Consider an independent two-side ranomdized design (Definition 2.1) where Assumption 2.1 holds. Let $\mathcal{C}^{(k)}$ be a $k$-block conditioning event as in Definition 3.1. Suppose also that Assumption 2.2 holds. Then, $\phi(\mathbf{W}^{obs};\mathcal{C}^{(k)})$ defined in Procedure 3.2 is finite-sample valid, such that $\mathbb{E}[\phi(\mathbf{W}^{obs};\mathcal{C}^{(k)})\mid\mathcal{C}^{(k)}]\leq\alpha$ for any finite $I,J$.
@@ 3.3.2 Selecting $k$ Based on Power Considerations
$\phi_{|\mathcal{U}|,|\mathcal{W}|}\geq\frac{1}{1+Ae^{-a\tau\sqrt{2nk}}}-O\left(C(2n/k,n/k)^{-0.5+\delta}\right)-\epsilon,$
@@
In Figure 3, we illustrate the lower bound on power as a function of the block size $k$, highlighting the trade-off between sample size and the number of unique treatment assignments as $k$ increases. While the plot identifies the optimal $k$ that maximizes power, in practice, the parameters of the lower bound function are unknown. Consequently, the optimal $k$ cannot be directly determined using Proposition 3.1. However, a heuristic approach can be used to select the largest feasible $k$ based on a predetermined maximum power threshold. Based on Proposition 3.1, $O\left(C(2n/k,n/k)^{-0.5+\delta}\right)$ controls the maximum power of the test, while $2nk$ determines how rapidly the power function reaches its maximum (sensitivity). Therefore, to optimize the detection of treatment effects, we recommend selecting the largest feasible block size $k$ based on a predetermined maximum power. For example, setting this power to 0.95 ensures the maximal utilization of observations to detect the minimal size of the treatment effect. In practice, we can choose $k$ to satisfy $C(2n/k,n/k)\approx\frac{1}{(1-\beta)^{2}}$, where $\beta$ is the predetermined power level. For example, consider the completely randomized two-sided design discussed before, where choosing a block size $k=n/6$ results in $|\mathcal{W}|=C(12,6)=924$, achieving a maximum power approximately equal to 0.967.
@@ 4 Randomization Tests for Weak Null Hypotheses
There has been a growing number of works on the study of randomization tests under weak null hypotheses; see, for example, (Chung and Romano, 2013; Li and Ding, 2017; Canay et al., 2017; DiCiccio and Romano, 2017a, b; Zhao and Ding, 2021; Wu and Ding, 2021). This literature has pointed out that randomization tests based on sharp nulls are not necessarily valid under the weak null. Nevertheless, in classic one-sided experiments, validity can often be restored by studentizing the test statistic. The most commonly used approach relies on a Neyman-style variance estimator, which coincides with the usual two-sample $t$-test variance.
In this section, we study use of conditional randomization tests with suitable test statistics and studentization for testing weak null hypotheses in two-sided randomization design. We document that the validity of these tests hinges on both the specific form of the null as well as the way in which studentization is carried out.
Specifically, we consider two types of weak nulls. The first, which we call the population average spillover effect from buyers, imposes a single mean-zero restriction on the average treatment effect after pooling over all buyers and sellers.
Hypothesis 4.1 (Population Average Spillover Effect from Buyers).
$H_{0}^{wb,1}:\frac{1}{IJ}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{i,j}(0,0)=\frac{1}{IJ}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{i,j}(1,0)~.$ (8)
The second, which we call the seller-specific average spillover effect from buyers, imposes a collection of mean-zero restrictions, one for each seller, requiring that the buyer-level average treatment effect vanish seller by seller.
Hypothesis 4.2 (Seller-specific Average Spillover Effect from Buyers).
$H_{0}^{wb,2}:\frac{1}{I}\sum_{i=1}^{I}Y_{i,j}(0,0)=\frac{1}{I}\sum_{i=1}^{I}Y_{i,j}(1,0)\quad\text{for all }j~.$ (9)
We show that the natural test statistic, together with commonly used Neyman-style studentization, leads to a valid test for the seller-specific null, but not for the population null. Therefore, we propose a modified, two-way (buyer $\times$ seller) variance estimator that restores asymptotic validity for the population-average weak null by absorbing the additional sampling variation induced by seller randomization.
4.1 Standard Neyman-style studentization
We begin by presenting the testing procedure under standard Neyman-style studentization:
Procedure 4.1 (Test Weak Null via standard (buyer-side) studentization).